OpenStudy (anonymous):

12x^(-1/2)+7x^(1/2)+x^3/2. Factor expression completely

jhonyy9 (jhonyy9):

12 ------ + 7radical x + radical x*3 = 12 radical x + 7 radical x + x radical x = x*(1/2) = (12 + 7 + x) radical x = (19 + x )radical x